# Global analysis of charm mixing parameters and determination of the CKM angle $\mathbf{\gamma}$

### v1

@@ Abstract
We present an updated global analysis of beauty decays sensitive to the angle $\gamma$ of the Cabibbo-Kobayashi-Maskawa matrix and of $D$-meson mixing data in the framework of approximate universality for charm mixing. We extract the fundamental theoretical parameters determining absorptive and dispersive contributions to $D$ meson mixing and CP violation, together with the angle $\gamma$. The results for the charm mixing parameters are $x_{12}\simeq x=(0.402\pm 0.034)\%$ and $y_{12}\simeq y=(0.629\pm 0.021)\%$, while the two CP-violating phases are given by $\phi_{2}^{M}=(2.4\pm 1.6)\degree$ and $\phi_{2}^{\Gamma}=(2.6\pm 1.2)\degree$. The angle $\gamma$ is found to be $\gamma=(68.4\pm 2.6)\degree$.
@@ I Introduction
Charm mixing occurs through Flavour Changing Neutral Currents (FCNC) that are absent at the tree-level in the Standard Model (SM) and suppressed by the Glashow-Iliopoulous-Maiani (GIM) mechanism and by the hierarchical structure of the Cabibbo-Kobayashi-Maskawa (CKM) matrix. Thus, heavy new physics (NP) may appear in the neutral $D$ system Grossman et al. (2007); Ciuchini et al. (2007); Golowich et al. (2007); Fajfer et al. (2007); Bona et al. (2008); Artuso et al. (2008); Brod et al. (2012a); Pirtskhalava and Uttayarat (2012); Bhattacharya et al. (2012); Cheng and Chiang (2012); Franco et al. (2012); Brod et al. (2012b); Giudice et al. (2012); Feldmann et al. (2012); Keren-Zur et al. (2013); Grossman and Schacht (2019); Dery and Nir (2019); Buras et al. (2021); Schacht and Soni (2022); Schacht (2023) through observables as the amount of CP violation. The latter could be used as an interesting benchmark of the SM since it is suppressed by the fourth power of the sine of the Cabibbo angle $\theta_{C}$.
@@
The experimental study of neutral $D$ mixing has been continuously progressing in the past two decades Bitenc et al. (2008); Aubert et al. (2009); del Amo Sanchez et al. (2010); Asner et al. (2012); Insler et al. (2012); Aaltonen et al. (2013); Peng et al. (2014); Malde et al. (2015); Aaij et al. (2016a, b, c, 2019a); Betti (2021); Ablikim et al. (2021); Aaij et al. (2021a); Ablikim et al. (2024); Pajero (2022); Zhang et al. (2006); Aubert et al. (2007); Aaij et al. (2017, 2024a); Aitala et al. (1999); Link et al. (2000); Csorna et al. (2002); Zupanc et al. (2009); Lees et al. (2013); Ablikim et al. (2015); Starič et al. (2016); Aaij et al. (2019b); Nayak et al. (2020); Aaij et al. (2022, 2023a); Ablikim et al. (2022); Aaij et al. (2019c, 2021b, 2023b, a, 2024a). The most precise estimates of the charm parameters are obtained from the rates of Cabibbo Favoured (CF) (i.e. proportional to $V_{cs}V_{ud}^{*}$) or Doubly Cabibbo Suppressed (DCS) (i.e. proportional to $V_{cd}V_{us}^{*}$) $D$ decays Zhang et al. (2006); Aubert et al. (2007); Aaij et al. (2017, 2024a), from charm decays to CP eigenstates Aitala et al. (1999); Link et al. (2000); Csorna et al. (2002); Zupanc et al. (2009); Lees et al. (2013); Ablikim et al. (2015); Starič et al. (2016); Aaij et al. (2019b); Nayak et al. (2020); Aaij et al. (2022, 2023a); Ablikim et al. (2022) and from Dalitz plot analyses of the three-body mode $K^{0}_{S}\pi^{+}\pi^{-}$ Aaij et al. (2019c, 2021b, 2023b). In particular, LHCb has achieved impressive results during the LHC Run II, including the observation of direct CP asymmetries in the Singly Cabibbo Suppressed (SCS) decays of $D$ mesons to $K^{+}K^{-}$ and $\pi^{+}\pi^{-}$ Aaij et al. (2019a, 2023a) and the very recent measurements of CP-violating observables in $D^{0}\to K^{\pm}\pi^{\mp}$ Aaij et al. (2024a).
@@
This remarkable increase in the experimental accuracy calls for an update of our previous works on $D$ meson mixing Bevan et al. (2012, 2014), generalizing the so-called “superweak assumption”, in which the weak phase describing CP violation in the interference between $D$ meson decays with and without absorptive mixing is set to zero, to the more suitable “approximate universality” scenario introduced in ref. Kagan and Silvestrini (2021) to be detailed below. To fully exploit the large amount of experimental information we include also beauty observables in the combination, allowing for a simultaneous determination of the CKM angle $\gamma=\arg(-V_{ud}V_{ub}^{*}V_{cb}V_{cd}^{*})$ from tree-level $B$ decays Abe et al. (2006); Horii et al. (2011); Nayak et al. (2013); Aaij et al. (2015a, b, 2016d); LHC (2016); Aaij et al. (2016e, 2019d, 2021c, 2020); Adachi et al. (2023, 2024a); Aaij et al. (2024b, c); Poluektov et al. (2010); Aaij et al. (2016f); Resmi et al. (2019); Aaij et al. (2021d); Abudinén et al. (2022); Aaij et al. (2023c, 2024d, d, 2024e, 2018, 2021e, 2024f) and of $D$ mixing parameters. Our results can be compared with the latest LHCb Aaij et al. (2021f); LHC (2024) estimates, the Belle/Belle II analysis Adachi et al. (2024b) for $\gamma$ and HFLAV Amhis et al. (2023) for the charm parameters. The consistency of the fit is checked by comparing $\gamma$ estimates obtained by combining measurements involving different types of $B$ mesons separately: charged $B^{\pm}$ or neutral $B^{0}$ and $B^{0}_{s}$.
@@
The paper is organized as follows: the dispersive-absorptive formalism for $D$ meson mixing in the framework of approximate universality is briefly summarized in Sec. II. The expressions of the charm observables are described in Sec. III, while the beauty observables are introduced in Sec. IV and their dependence on the parameters of interest of this study are reported in Apps. A and B. We present and discuss the most relevant results in Sec. V. A complete list of the parameters introduced and of their prior distributions is reported in App. C, while their posterior probability intervals are given in App. D. A summary can be found in Sec. VI.
@@ II Charm Mixing Formalism
The time evolution of a linear combination of $D^{0}$ and $\overline{D^{0}}$ mesons follows the Schrödinger equation, with a $2\times 2$ non-Hermitian Hamiltonian $\mathbf{H}$ (see e.g. Nir (1992); Branco et al. (1999); Grossman et al. (2009); Kagan and Sokoloff (2009)) that in terms of its dispersive $\mathbf{M}$ and absorptive $\mathbf{\Gamma}$ components is given by
@@
In the SM, it is possible to decompose the dispersive and absorptive parts of $H_{12}$ in terms of U-spin amplitudes as follows Kagan and Silvestrini (2021):
@@
Employing in the definition of Eq. (9) the expressions of $M_{12}^{\text{SM}}$ and $\Gamma_{12}^{\text{SM}}$ given in Eq. (8) and neglecting the third (smallest) contribution, the orders of magnitude of the two CP-violating phases in the SM can be roughly estimated to be Kagan and Silvestrini (2021)
@@
From Eqs. (6) and (8), we get the relation $\left|\Gamma_{2}\right|\simeq y_{12}\ \Gamma/\left|\lambda^{s}_{uc}\right|^{2}$, which can be employed in Eq. (10) to provide an upper bound on $\left|\phi_{2}^{\Gamma}\right|$ as Kagan and Silvestrini (2021)
$\left|\phi_{2}^{\Gamma}\right|\simeq\frac{\left|\lambda^{b}_{uc}\lambda^{s}_{uc}\sin\gamma\right|}{y_{12}}\times\frac{\left|\Gamma_{1}\right|}{\Gamma}<(5\times 10^{-3})\times\bigg{(}\frac{0.63\%}{y_{12}}\bigg{)},$ (11)
@@
Notice that $\phi_{2}$ corresponds to the phase $\phi$ extracted in global analyses under the assumption of final state independence (see e.g. refs. Aaij et al. (2021f); LHC (2024)).
In the superweak assumption, dispersive mixing is supposed to be the dominant source of CPV in the charm system, corresponding to the approximation where $\phi_{2}^{\Gamma}=0$ and $\phi_{12}=\phi_{2}^{M}$ or, equivalently, $\phi=\phi_{2}=\arctan\frac{1-|q/p|^{2}}{1+|q/p|^{2}}\frac{x}{y}$. Then, going from the superweak assumption, adopted in our previous works Bevan et al. (2012, 2014), to approximate universality, used in this work, involves moving from a description in terms of a single independent CPV parameter ($\phi_{12}$ or $\phi_{2}$) to a scenario with two free CPV parameters ($\phi_{2}^{M}$ and $\phi_{2}^{\Gamma}$ or $\phi_{2}$ and $|q/p|-1$).
@@ III Charm observables
The charm observables used in our combination rely on time-integrated and time-dependent studies of $D$ meson decays, reconstructed from two- and multi-body final states. The final states can be CP eigenstates or originated from CF or DCS decays. We parametrize the decay amplitudes through the ratios of their magnitudes and strong phases and provide expressions for the observables in Tab. 1 in terms of charm decay, mixing and CP-violating parameters in the framework of approximate universality for the neutral $D$ system.
@@ III.1 CF/DCS decays to $\mathbf{K^{\pm}X}$
The observables used in the fit for this type of final states are extracted from the ratios of the WS and RS decay rates Aaij et al. (2017); Aubert et al. (2007); Zhang et al. (2006), given by
@@
The linear and quadratic terms in Eq. (15) can be decomposed into
$x^{\prime\pm}_{K\pi\pi^{0}}$, $y^{\prime\pm}_{K\pi\pi^{0}}$ $D\to K\pi\pi^{0}$ Aubert et al. (2009) $(x^{2}+y^{2})/2$ $D\to Kl\overline{\nu}_{l}$ Bitenc et al. (2008)
$\cos\delta_{D}^{K\pi}$, $D\to K\pi$ Asner et al. (2012) $\delta_{D}^{K^{0}_{S}K\pi}$, $D\to K^{0}_{S}K\pi$ Insler et al. (2012)
$\sin\delta_{D}^{K\pi}$ $\kappa_{D}^{K^{0}_{S}K\pi}$
$\tilde{y}_{\text{CP}}$ $D\to XX$ Aitala et al. (1999); Link et al. (2000); Csorna et al. (2002); Zupanc et al. (2009); Lees et al. (2013); Ablikim et al. (2015); Starič et al. (2016); Aaij et al. (2019b); Nayak et al. (2020); Aaij et al. (2022) $x$, $y$, $|q/p|$, $\phi_{2}$ $D\to K^{0}_{S}\pi\pi$ Peng et al. (2014)
$F_{+}^{XX\pi^{0}}$ $D\to XX\pi^{0}$ Malde et al. (2015) $x$, $y$ $D\to K^{0}_{S}\pi\pi$ del Amo Sanchez et al. (2010); Aaij et al. (2016a)
$F_{+}^{4\pi}$ $D\to 4\pi$ Malde et al. (2015); Ablikim et al. (2024) $\frac{\mathcal{B}(D\to K^{0}_{s}K^{+}\pi^{-})}{\mathcal{B}(D^{0}\to K^{0}_{s}K^{-}\pi^{+})}$ $D\to K^{0}_{S}K\pi$ Aaij et al. (2016b)
$(r_{D}^{K\pi})^{2}$, $(x^{\prime\pm}_{K\pi})^{2}$, $y^{\prime\pm}_{K\pi}$ $D\to K\pi$ Zhang et al. (2006); Aubert et al. (2007); Aaij et al. (2017) $(x^{2}+y^{2})/4$ $D\to K3\pi$ Aaij et al. (2016c)
$x_{\text{CP}}$, $y_{\text{CP}}$, $\Delta x$, $\Delta y$ $D\to K^{0}_{S}\pi\pi$ Aaij et al. (2019c, 2021b, 2023b) $\Delta A_{D}^{\text{CP}}$, $(\left\langle\tau\right\rangle^{KK}_{E^{\prime}}\pm\left\langle\tau\right\rangle^{\pi\pi}_{E^{\prime}})$ $D\to XX$ Aaij et al. (2019a)
$r_{D}^{K3\pi}$, $\delta_{D}^{K3\pi}$, $\kappa_{D}^{K3\pi}$, $D\to K3\pi$ Ablikim et al. (2021) $\Delta Y^{KK}-\Delta Y^{\pi\pi}$ $D\to XX$ Aaij et al. (2021a)
$r_{D}^{K\pi\pi^{0}}$, $\delta_{D}^{K\pi\pi^{0}}$, $\kappa_{D}^{K\pi\pi^{0}}$ $D\to K\pi\pi^{0}$ $\Delta Y$
$(r_{D}^{K\pi})^{2}$, $c^{(^{\prime})}_{K\pi}$, $\Delta c^{(^{\prime})}_{K\pi}$, $D\to K\pi$ Aaij et al. (2024a) $A_{D}(f^{+})$, $A_{D}(\pi\pi\pi^{0})$, $D\to f^{+}$, $D\to K\pi$ Ablikim et al. (2022)
$\Delta\tilde{c}^{(^{\prime})}_{K\pi}$ $r_{D}^{K\pi}\cos\delta_{D}^{K\pi}$, $r_{D}^{K\pi}\sin\delta_{D}^{K\pi}$ $D\to XX\pi^{0}$
$A_{D}^{\text{CP}}(KK)$, $\left\langle\tau\right\rangle_{E}^{KK}$ $D\to K^{+}K^{-}$ Aaij et al. (2023a)
their CP-conserving ($c_{f}$, $c^{\prime}_{f}$) and CP-violating parts ($\Delta c_{f}$, $\Delta c^{\prime}_{f}$), as follows:
@@
The separation of the CP-violating from the CP-conserving contributions entering the WS/RS ratios of $D^{0}\to K^{\pm}\pi^{\mp}$ decays has been performed for the first time this year by LHCb Aaij et al. (2024a), combining both the Run 1 and Run 2 results. However, while the Run 1 measurements can be used directly to extract the coefficients in Eq. (19), the Run 2 data use experimental information from $D^{0}\to K^{+}K^{-}$ decays to correct a detection asymmetry. This operation leaves invariant $c_{f}$ and $c^{\prime}_{f}$, while it changes the expressions of the CPV observables as
@@
and the parameters $r_{D}^{f}$ are replaced by $r_{D}^{f}(1\mp a_{D}^{KK})$ in Eq. (15). Here, $a_{D}^{KK}$ is the direct CP asymmetry in $D^{0}\to K^{+}K^{-}$ decays and we introduced the parameter $\Delta Y^{KK}$. They were measured in refs. Aaij et al. (2021a) and Aaij et al. (2023a), and are defined in terms of decay amplitudes and decay rates in Eqs. (23) and (26), respectively, when describing the observables involving SCS decays to CP eigenstates.
We consider in the combination also observables coming from the analysis of quantum correlated $D^{0}-\overline{D^{0}}$ pairs (see e.g. Ablikim et al. (2022)). Here, one of the $D$ mesons decays to the so-called tagging mode, while the other decays to the signal mode, typically $K^{-}\pi^{+}$. Then, the following asymmetries are measured:

### v3

@@ Abstract
We present an updated global analysis of beauty decays sensitive to the angle $\gamma$ of the Cabibbo-Kobayashi-Maskawa matrix and of $D$-meson mixing data in the framework of approximate universality, in which CP violation in $D-\overline{D}$ mixing is described in terms of two universal weak phases corresponding to dispersive and absorptive contributions. We extract the fundamental theoretical parameters determining absorptive and dispersive contributions to $D$ meson mixing and CP violation, together with the angle $\gamma$. The results for the charm mixing parameters are $x_{12}\simeq x=(0.401\pm 0.043)\%$ and $y_{12}\simeq y=(0.610\pm 0.017)\%$, while the two CP-violating phases are given by $\phi_{2}^{M}=(0.13\pm 0.70)\degree$ and $\phi_{2}^{\Gamma}=(2.1\pm 1.6)\degree$. The angle $\gamma$ is found to be $\gamma=(65.7\pm 2.5)\degree$, in excellent agreement with the indirect determination from the Unitarity Triangle analysis.
@@ I Introduction
Charm mixing occurs through Flavour Changing Neutral Currents (FCNC) that are absent at the tree-level in the Standard Model (SM) and suppressed by the Glashow-Iliopoulous-Maiani (GIM) mechanism Glashow:1970gm and by the hierarchical structure of the Cabibbo-Kobayashi-Maskawa (CKM) matrix Cabibbo:1963yz ; Kobayashi:1973fv . Thus, new physics (NP) effects due to heavy particles may appear in the neutral $D$ system Grossman:2006jg ; Ciuchini:2007cw ; Golowich:2007ka ; Fajfer:2007dy ; UTfit:2007eik ; Artuso:2008vf ; Brod:2011re ; Pirtskhalava:2011va ; Bhattacharya:2012ah ; Cheng:2012wr ; Franco:2012ck ; Brod:2012ud ; Giudice:2012qq ; Feldmann:2012js ; Keren-Zur:2012buf ; Grossman:2019xcj ; Dery:2019ysp ; Buras:2021rdg ; Schacht:2021jaz ; Schacht:2022kuj through observables such as the amount of CP violation. The latter could be used as an interesting benchmark of the SM since it is suppressed by the fourth power of the sine of the Cabibbo angle $\theta_{C}$.
@@
The experimental study of neutral $D$ mixing has been continuously progressing in the past two decades E791:1996klq ; CLEO:2005oam ; BaBar:2004grg ; BaBar:2007fup ; Belle:2008qhk ; Belle:2008qhk ; BaBar:2008xkf ; BaBar:2010nhz ; LHCb:2015lgi ; BaBar:2016kvp ; CLEO:2012fel ; CLEO:2012obf ; CDF:2013gvz ; Belle:2014ydf ; LHCb:2019mxy ; LHCb:2022cak ; Malde:2015mha ; BESIII:2024nnf ; BESIII:2024zco ; LHCb:2015lnk ; BESIII:2025vbt ; BESIII:2025ypr ; LHCb:2016zmn ; Betti:2021mpf ; Libby:2014rea ; Evans:2016tlp ; BESIII:2021eud ; BESIII:2022qkh ; Pajero:2022vev ; Belle:2006ipk ; BaBar:2007kib ; LHCb:2025kch ; LHCb:2024hyb ; E791:1999bzz ; FOCUS:2000kxx ; CLEO:2001lgl ; Belle:2009xzl ; BaBar:2012bho ; BESIII:2015ado ; Belle:2015etc ; LHCb:2018zpj ; Belle:2019xha ; LHCb:2022gnc ; CDF:2014wyb ; LHCb:2015xyd ; LHCb:2017ejh ; LHCb:2019dom ; LHCb:2021vmn ; LHCb:2016nxk ; LHCb:2022lry ; LHCb:2014kcb ; LHCb:2016csn ; LHCb:2019hro ; BaBar:2007tfw ; Belle:2008ddg ; DiCanto:2012ufu ; CDF:2011ejf ; LHCb:2024hyb . The most precise estimates of the charm parameters are obtained from the rates of Cabibbo Favoured (CF) (i.e. proportional to $V_{cs}V_{ud}^{*}$) or Doubly Cabibbo Suppressed (DCS) (i.e. proportional to $V_{cd}V_{us}^{*}$) $D$ decays Belle:2006ipk ; BaBar:2007kib ; LHCb:2025kch ; LHCb:2024hyb , from charm decays to CP eigenstates E791:1999bzz ; FOCUS:2000kxx ; CLEO:2001lgl ; Belle:2009xzl ; BaBar:2012bho ; BESIII:2015ado ; Belle:2015etc ; LHCb:2018zpj ; Belle:2019xha ; LHCb:2022gnc ; CDF:2014wyb ; LHCb:2015xyd ; LHCb:2017ejh ; LHCb:2019dom ; LHCb:2021vmn ; LHCb:2016nxk ; LHCb:2022lry ; LHCb:2014kcb ; LHCb:2016csn ; LHCb:2019hro ; BaBar:2007tfw ; Belle:2008ddg ; DiCanto:2012ufu ; CDF:2011ejf and from Dalitz plot analyses of the three-body modes $K^{0}_{S}\pi^{+}\pi^{-}$, $K^{0}_{S}K^{+}K^{-}$ or $\pi^{+}\pi^{-}\pi^{0}$ BaBar:2010nhz ; LHCb:2015lgi ; Belle:2014ydf ; BaBar:2016kvp ; LHCb:2019mxy ; LHCb:2022cak . In particular, LHCb has achieved impressive results during the LHC Run II, including the observation of direct CP asymmetries in the Singly Cabibbo Suppressed (SCS) decays of $D$ mesons to $K^{+}K^{-}$ and $\pi^{+}\pi^{-}$ LHCb:2019hro ; LHCb:2022lry and the very recent measurements of CP-violating observables in $D^{0}\to K^{\pm}\pi^{\mp}$ LHCb:2024hyb .
@@
This remarkable increase in the experimental accuracy calls for an update of our previous works on $D$ meson mixing UTfit:2012ich ; UTfit:2014hez , generalizing the so-called “superweak assumption”, in which the weak phase describing CP violation in the interference between $D$ meson decays with and without absorptive mixing is set to zero, to the more suitable “approximate universality” scenario introduced in Ref. Kagan:2020vri to be detailed below.
Charm mixing also plays an important role in the extraction of the CKM angle $\gamma=\arg(-V_{ud}V_{ub}^{*}V_{cb}V_{cd}^{*})$ from tree-level $B$ decays Belle:2006cuz ; Belle:2011ac ; Belle:2013dtr ; LHCb:2015dlc ; LHCb:2017egy ; LHCb:2016bsl ; LHCb:2020hdx ; LHCb:2020vut ; LHCb:2021mmv ; Belle:2023lha ; Belle:2023yoe ; LHCb:2024oco ; LHCb:2022nng ; LHCb:2024ett ; Belle:2010xyn ; Belle:2019uav ; LHCb:2020yot ; Belle:2021efh ; LHCb:2023yjo ; LHCb:2023lib ; LHCb:2023kpr ; LHCb:2023ayf ; LHCb:2018zap ; LHCb:2020qag ; LHCB-PAPER-2024-020 ; BaBar:2005jis ; BaBar:2006slj ; Belle:2006lts ; Belle:2011dhx ; BaBar:2007dro ; BaBar:2008qcq ; BaBar:2009dzx ; BaBar:2010hvw ; BaBar:2010otv ; BaBar:2011rud ; CDF:2009wnr ; CDF:2011xrp ; Belle:2006lys ; BaBar:2010uep ; Belle:2015roy , calling for a simultaneous combination of charm and beauty measurements, as pioneered by the LHCb collaboration LHCb:2021dcr . Our results can be compared with the latest LHCb LHCb-CONF-2024-004 estimates, the Belle/Belle II analysis Belle-II:2024eob for $\gamma$ and HFLAV HFLAV:2022pwe for the charm parameters. The consistency of the fit is checked by comparing $\gamma$ estimates obtained by combining measurements involving different types of $B$ mesons separately: charged $B^{\pm}$ or neutral $B^{0}$ and $B^{0}_{s}$. The impact of the global fit on the charm parameters is assessed by comparing it with the results of a fit to $D$ observables only.
@@
The paper is organized as follows: the dispersive-absorptive formalism for $D$ meson mixing in the framework of approximate universality is briefly summarized in Sec. II. The expressions of the charm observables are described in Sec. III, while the beauty observables are introduced in Sec. IV and their dependence on the parameters of interest of this study is reported in Apps. A and B. We present and discuss the most relevant results in Sec. V. The posterior probability intervals of the fit parameters are given in App. C. A summary can be found in Sec. VI.
@@ II Charm Mixing Formalism
The time evolution of a linear combination of $D^{0}$ and $\overline{D^{0}}$ mesons follows the Schrödinger equation, with a $2\times 2$ non-Hermitian Hamiltonian $\mathbf{H}$ (see e.g. Nir:1992uv ; Branco:1999fs ; Grossman:2009mn ; Kagan:2009gb ) that in terms of its dispersive $\mathbf{M}$ and absorptive $\mathbf{\Gamma}$ components is given by
@@
In the SM, it is possible to decompose the dispersive and absorptive parts of $H_{12}$ in terms of U-spin amplitudes as follows Kagan:2020vri :
@@
Employing in the definition of Eq. (9) the expressions of $M_{12}^{\text{SM}}$ and $\Gamma_{12}^{\text{SM}}$ given in Eq. (8) and neglecting the third (smallest) contribution, the orders of magnitude of the two CP-violating phases in the SM can be roughly estimated to be Kagan:2020vri
@@
From Eqs. (6) and (8), we get the relation $\left|\Gamma_{2}\right|\simeq y_{12}\ \Gamma/\left|\lambda^{s}_{uc}\right|^{2}$, which can be employed in Eq. (10) to provide an upper bound on $\left|\phi_{2}^{\Gamma}\right|$ as Kagan:2020vri
$\left|\phi_{2}^{\Gamma}\right|\simeq\frac{\left|\lambda^{b}_{uc}\lambda^{s}_{uc}\sin\gamma\right|}{y_{12}}\times\frac{\left|\Gamma_{1}\right|}{\Gamma}<(5\times 10^{-3})\times\bigg{(}\frac{0.61\%}{y_{12}}\bigg{)},$ (11)
@@
Notice that $\phi_{2}$ corresponds to the phase $\phi$ extracted in global analyses under the assumption of final state independence (see e.g. refs. LHCb:2021dcr ; LHCb-CONF-2024-004 ).
In the superweak assumption, dispersive mixing is supposed to be the dominant source of CPV in the charm system, corresponding to the approximation where $\phi_{2}^{\Gamma}=0$ and $\phi_{12}=\phi_{2}^{M}$ or, equivalently, $\phi=\phi_{2}=\arctan\frac{1-|q/p|^{2}}{1+|q/p|^{2}}\frac{x}{y}$. Then, going from the superweak assumption, adopted in our previous works UTfit:2012ich ; UTfit:2014hez , to approximate universality, used in this work, involves moving from a description in terms of a single independent CPV parameter ($\phi_{12}$ or $\phi_{2}$) to a scenario with two free CPV parameters ($\phi_{2}^{M}$ and $\phi_{2}^{\Gamma}$ or $\phi_{2}$ and $|q/p|-1$).
@@ III Charm observables
The charm observables used in our combination rely on time-integrated and time-dependent studies of $D$ meson decays, reconstructed from two- and multi-body final states. The final states can be either CP eigenstates or non CP eigenstates originated from CF or DCS decays. We parametrize the decay amplitudes through the ratios of their magnitudes and strong phases and provide expressions for the observables in Tab. 1 in terms of charm decay, mixing and CP-violating parameters in the framework of approximate universality for the neutral $D$ system, following Ref. Kagan:2020vri .
@@ III.1 CF/DCS decays to $\mathbf{K^{\pm}X}$
$(x^{2}+y^{2})/2$ $D\to Kl\overline{\nu}_{l}$ E791:1996klq ; CLEO:2005oam ; BaBar:2004grg ; BaBar:2007fup ; Belle:2008qhk $x^{\prime\pm}_{K\pi\pi^{0}}$, $y^{\prime\pm}_{K\pi\pi^{0}}$ $D\to K\pi\pi^{0}$ BaBar:2008xkf
$x$, $y$ $D\to K^{0}_{S}XX$ BaBar:2010nhz $x$, $y$ $D\to K^{0}_{S}\pi\pi$ LHCb:2015lgi
$x$, $y$ $D\to\pi^{+}\pi^{-}\pi^{0}$ BaBar:2016kvp $x$, $y$, $|q/p|$, $\phi_{2}$ $D\to K^{0}_{S}\pi\pi$ Belle:2014ydf
$\cos\delta_{D}^{K\pi}$, $D\to K\pi$ CLEO:2012fel $\delta_{D}^{K^{0}_{S}K\pi}$, $D\to K^{0}_{S}K\pi$ CLEO:2012obf
$\sin\delta_{D}^{K\pi}$ $\kappa_{D}^{K^{0}_{S}K\pi}$
$F_{+}^{4\pi}$ $D\to 4\pi$ Malde:2015mha ; BESIII:2024zco $F_{+}^{XX\pi^{0}}$ $D\to XX\pi^{0}$ Malde:2015mha ; BESIII:2024nnf
$\frac{\mathcal{B}(D\to K^{0}_{s}K^{+}\pi^{-})}{\mathcal{B}(D^{0}\to K^{0}_{s}K^{-}\pi^{+})}$ $D\to K^{0}_{S}K\pi$ LHCb:2015lnk $F_{+}^{KK\pi\pi}$ $D\to K^{+}K^{-}\pi^{+}\pi^{-}$ BESIII:2025ypr
$x_{\text{CP}}$, $y_{\text{CP}}$, $\Delta x$, $\Delta y$ $D\to K^{0}_{S}\pi\pi$ LHCb:2019mxy ; LHCb:2022cak $(x^{2}+y^{2})/4$ $D\to K3\pi$ LHCb:2016zmn
$r_{D}^{K3\pi}$, $\delta_{D}^{K3\pi}$, $\kappa_{D}^{K3\pi}$, $D\to K3\pi$ Libby:2014rea ; Evans:2016tlp ; LHCb:2016zmn ; BESIII:2021eud $A_{D}(f^{+})$, $A_{D}(\pi\pi\pi^{0})$, $D\to f^{+}$, $D\to K\pi$ BESIII:2022qkh
$r_{D}^{K\pi\pi^{0}}$, $\delta_{D}^{K\pi\pi^{0}}$, $\kappa_{D}^{K\pi\pi^{0}}$ $D\to K\pi\pi^{0}$ $r_{D}^{K\pi}\cos\delta_{D}^{K\pi}$, $r_{D}^{K\pi}\sin\delta_{D}^{K\pi}$ $D\to XX\pi^{0}$
$\tilde{y}_{\text{CP}}$ $D\to XX$ E791:1999bzz ; FOCUS:2000kxx ; CLEO:2001lgl ; Belle:2009xzl ; BaBar:2012bho ; BESIII:2015ado ; Belle:2015etc ; LHCb:2018zpj ; Belle:2019xha ; LHCb:2022gnc $\Delta Y^{KK}-\Delta Y^{\pi\pi}$, $\Delta Y$ $D\to XX$ LHCb:2015xyd ; LHCb:2017ejh ; LHCb:2019dom ; LHCb:2021vmn
$(r_{D}^{K\pi})^{2}$, $c^{(^{\prime})}_{K\pi}$, $\Delta c^{(^{\prime})}_{K\pi}$, $\Delta\tilde{c}^{(^{\prime})}_{K\pi}$ $D\to K\pi$ LHCb:2024hyb $(r_{D}^{K\pi})^{2}$, $(x^{\prime\pm}_{K\pi})^{2}$, $y^{\prime\pm}_{K\pi}$ $D\to K\pi$ Belle:2006ipk ; BaBar:2007kib ; LHCb:2025kch
$A_{D}^{\text{CP}}(KK)$, $\left\langle\tau\right\rangle_{E}^{KK}$ $D\to K^{+}K^{-}$ LHCb:2014kcb ; LHCb:2016nxk ; LHCb:2022lry $\Delta A_{D}^{\text{CP}}$, $\left\langle\tau\right\rangle^{KK}_{E}$, $\left\langle\tau\right\rangle^{\pi\pi}_{E}$ $D\to XX$ LHCb:2014kcb ; LHCb:2016csn ; LHCb:2019hro
$\Delta Y^{KK}$, $\Delta Y^{\pi\pi}$ $D\to XX$ CDF:2014wyb $A_{D}^{\text{CP}}(XX)$, $\left\langle\tau\right\rangle_{E}^{KK}$, $\left\langle\tau\right\rangle_{E}^{\pi\pi}$ $D\to XX$ BaBar:2007tfw ; Belle:2008ddg ; DiCanto:2012ufu ; CDF:2011ejf
$\frac{\mathcal{B}(D^{0}\to K^{+}\pi^{-})}{\mathcal{B}(D^{0}\to K^{-}\pi^{+})}$, $\frac{\mathcal{B}(D^{0}\to K^{+}\pi^{-}\pi^{0})}{\mathcal{B}(D^{0}\to K^{-}\pi^{+}\pi^{0})}$ $D\to K\pi$, $D\to K\pi\pi^{0}$ BESIII:2025vbt
$\frac{\mathcal{B}(D^{0}\to K^{+}\pi^{-}\pi^{+}\pi^{-})}{\mathcal{B}(D^{0}\to K^{-}\pi^{+}\pi^{+}\pi^{-})}$ $D\to K3\pi$
@@
The observables used in the fit for this type of final states are extracted from the ratios of the WS and RS decay rates LHCb:2025kch ; BaBar:2007kib ; Belle:2006ipk , given by
@@
The linear and quadratic terms in Eq. (15) can be decomposed into their CP-conserving ($c_{f}$, $c^{\prime}_{f}$) and CP-violating parts ($\Delta c_{f}$, $\Delta c^{\prime}_{f}$), as follows:
@@
The separation of the CP-violating from the CP-conserving contributions entering the WS/RS ratios of $D^{0}\to K^{\pm}\pi^{\mp}$ decays has been performed for the first time this year by LHCb LHCb:2024hyb , combining both the Run 1 and Run 2 results. However, while the Run 1 measurements can be used directly to extract the coefficients in Eq. (19), the Run 2 data use experimental information from $D^{0}\to K^{+}K^{-}$ decays to correct a detection asymmetry. This operation leaves invariant $c_{f}$ and $c^{\prime}_{f}$, while it changes the expressions of the CPV observables as
@@
and the parameters $r_{D}^{f}$ are replaced by $r_{D}^{f}(1\mp a_{D}^{KK})$ in Eq. (15). Here, $a_{D}^{KK}$ is the direct CP asymmetry in $D^{0}\to K^{+}K^{-}$ decays and we introduced the parameter $\Delta Y^{KK}$. They were measured in refs. LHCb:2014kcb ; LHCb:2016nxk ; LHCb:2022lry ; BaBar:2007tfw ; Belle:2008ddg ; DiCanto:2012ufu ; CDF:2011ejf and CDF:2014wyb ; LHCb:2015xyd ; LHCb:2017ejh ; LHCb:2019dom ; LHCb:2021vmn , and are defined in terms of decay amplitudes and decay rates in Eqs. (23) and (26), respectively, when describing the observables involving SCS decays to CP eigenstates.
We consider in the combination also observables coming from the analysis of quantum correlated $D^{0}-\overline{D^{0}}$ pairs (see e.g. BESIII:2022qkh ). Here, one of the $D$ mesons decays to the so-called tagging mode, while the other decays to the signal mode, typically $K^{-}\pi^{+}$. Then, the following asymmetries are measured: